OpenStudy (anonymous):

What is the correct classification of the system of equations below? 2x + 6y = 8 y - 2x = 2

OpenStudy (anonymous):

@shelbygt520

OpenStudy (texaschic101):

what do you mean classification ? There is 1 solution to this problem....is that what you mean ?

OpenStudy (anonymous):

parallel coincident intersecting

OpenStudy (texaschic101):

intersecting

OpenStudy (anonymous):

Intersecting.

OpenStudy (anonymous):

ok thx

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

What is the correct classification of the system of equations below? 12x - 3y = 9 y - 4x = -3 parallel coincident intersecting

OpenStudy (texaschic101):

12x - 3y = 9 -- reduce 4x - y = 3 4x - 3 = y y - 4x = -3 y = 4x - 3 same line coincident

OpenStudy (anonymous):

I'm just going to go ahead and let @texaschic101 take this down she's got it pretty good lol

OpenStudy (texaschic101):

thank you shelby :)

OpenStudy (anonymous):

thx :)

OpenStudy (texaschic101):

you got any more questions ?

OpenStudy (anonymous):

14x + 2y = 10 y + 7x = -5

OpenStudy (texaschic101):

14x + 2y = 10 -- reduce 7x + y = 5 y = -7x + 5 y + 7x = -5 y = -7x - 5 same slope, different y intercept parallel

OpenStudy (anonymous):

y - 3x = 3 y = 3x - 2

OpenStudy (texaschic101):

to do these problems, you need to put the equations in y = mx + b form, and compare the slopes(m). If they have same slope and different y intercepts, they are parallel. If they have different slopes, then there is 1 solution. If they have same slope and same y intercepts, they are coincident. y - 3x = 3 y = 3x + 3 y = 3x - 2 same slope, different y intercept parallel

OpenStudy (anonymous):

5x + 6y = 1 y = -5x + 6

OpenStudy (texaschic101):

5x + 6y = 1 6y = -5x + 1 y = -5/6x + 1/6 y = -5x + 6 one solution intersecting

OpenStudy (anonymous):

A bag contains 2 steel balls and 5 brass balls. The total weight is 13 pounds. If 2 steel balls are added and 2 brass balls are removed, the weight decreases to 12 pounds. How much does each kind of ball weigh? steel, 4 pounds; brass, 1 pound steel, 1.5 pounds; brass, 2 pounds steel, 0.75 pounds; brass, 3 pounds steel, 3 pounds; brass, 4 pounds

OpenStudy (texaschic101):

2s + 5b = 13 -->(-2) 4s + 3b = 12 ------------ -4s -10b = -26 4s + 3b = 12 ------------ -7b = -14 b = 2 now sub 2 in for b 2s + 5b = 13 2s + 5(2) = 13 2s + 10 = 13 2s = 13 - 10 2s = 3 s = 3/2 or 1 1/2 or 1.5 answer is B

OpenStudy (anonymous):

thx all I need is the answer :D yoru the best!!!!!!

OpenStudy (anonymous):

James purchased 2 T-shirts and a pair of jeans online and paid $80 for his purchase. The next week, the same site was offering a 50% discount on all its clothes. James bought 3 T-shirts and two pairs of jeans for $70. What were the original prices of T-shirts and jeans? (Assume that all T-shirts had the same original price, as did all pairs of jeans.) T-shirt, $25; jeans, $30 T-shirt, $20; jeans, $30 T-shirt, $20; jeans, $40 T-shirt, $25; jeans, $35

OpenStudy (texaschic101):

I am having trouble with this one...I am not getting any of your answer choices

OpenStudy (anonymous):

Given equation A as 2x+1/4y=3 and equation B as 2/3x-y=6 , which expression will give the value of y? A − 3B A − 3/2 B 2A − 3B 1/3 A + B

OpenStudy (anonymous):

just put answer

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

I do not know this...sorry. I have to go now....good luck

OpenStudy (anonymous):

@shelbygt520

OpenStudy (anonymous):

Given equation A as 2x+1/4y=3 and equation B as 2/3x-y=6 , which expression will give the value of y? A − 3B A − 3/2 B 2A − 3B 1/3 A + B